OpenStudy (anonymous):

Show that \[ \sum _{n=1}^{\infty } \frac{(2 n-1)!}{2^{4 n} (n!)^2}=\ln \left(-4 \left(\sqrt{3}-2\right)\right) \]

OpenStudy (anonymous):

Knowing the above property, one can give a nice solution to http://openstudy.com/study#/updates/54a2cb99e4b054f0c3b448b0 @ganeshie8 @cambrige @SithsAndGiggles

OpenStudy (anonymous):

Is there any chance we could write the LHS as a power series of a known function with \(x=\dfrac{1}{2}\)?

OpenStudy (anonymous):

Or \(x=\dfrac{1}{16}\)?

OpenStudy (kainui):

Just at first glance I recognized that we can rewrite some of this with even and odd factorial functions but after making some soup I don't know what else I can really do.

OpenStudy (anonymous):

Soup strikes again...

OpenStudy (kainui):

My only other strategy is turn lnx into a power series and then play around with the terms and try to make them look alike but... It doesn't look very promising lol.

OpenStudy (anonymous):

It somewhat resembles the Bessel function of the first kind: http://en.wikipedia.org/wiki/Bessel_function#Bessel_functions_of_the_first_kind_:_J.CE.B1 if it weren't for the \((2n-1)!\) in the numerator.

OpenStudy (anonymous):

Or rather the modified version: http://en.wikipedia.org/wiki/Bessel_function#Modified_Bessel_functions_:_I.CE.B1.2C_K.CE.B1

OpenStudy (kainui):

I've been studying tensors all morning and I'm burnt out on this kind of hard stuff haha. I'm gonna try to help some kids with slope intercept lol

OpenStudy (anonymous):

Hyperbolic trig perhaps?

OpenStudy (anonymous):

We have a winner: http://en.wikipedia.org/wiki/Inverse_hyperbolic_function#Series_expansions

OpenStudy (kainui):

Wow, that makes the natural logarithm part make sense quite a bit, now that I see that, I should have thought that earlier, good call.

OpenStudy (anonymous):

So if \[\newcommand{\arcsech}{\text{arcsech}} \arcsech x=\ln\frac{2}{x}-\sum_{n=1}^\infty \frac{(2n-1)!}{2^{2n}(n!)^2}x^{2n}\] then \[\arcsech \frac{1}{2}=\ln4-\sum_{n=1}^\infty \frac{(2n-1)!}{2^{4n}(n!)^2}~~\iff~~\sum_{n=1}^\infty \frac{(2n-1)!}{2^{2n}(n!)^2}x^{2n}=\ln4-\ln(2+\sqrt3)\]

OpenStudy (anonymous):

... and some minor algebraic manipulation seals the deal! Great question!

OpenStudy (anonymous):

\[\arcsech \frac{1}{2}=\ln4-\sum_{n=1}^\infty \frac{(2n-1)!}{2^{4n}(n!)^2}~~\iff~~\sum_{n=1}^\infty \frac{(2n-1)!}{2^{\color{red}{4n}}(n!)^2}=\cdots\]